solve the linear system...if the system is inconsistent or dependent, so state
3x-2y = 0
9x-8y= 7

Answer by checkley71(8403)

3X-2Y=0 OR X=2Y/3 SUBSTITUTING 2Y/3 FOR X IN EQUATION 2 WE GET 9*2Y/3-8Y=7 OR
18Y/3-8Y=7 OR 6Y-8Y=7 OR -2Y=7 OR Y=-7/2 THUS X=2(-7/2)/3 X=-7/3 OR
3*(-7/3)-2(-7/2)=0 OR -7+7=0 OR 0=0